OpenStudy (anonymous):

use the intercepts to graph the equation x-5=y

OpenStudy (saifoo.khan):

whn y intercpts x = 0 whn x intercpets y = 0

OpenStudy (aroub):

x-intercept = (0,0) y-intercept=(0,-5)

OpenStudy (saifoo.khan):

so x intercpt equals x - 5 = 0 x = 5 ================================= y interxpt equals 0 - 5 = y y = 5

OpenStudy (aroub):

x-intercept=(5,0) *****

OpenStudy (saifoo.khan):

x-intercept = (5,0) y-intercept=(0,-5)

OpenStudy (saifoo.khan):

as aroub said.